OpenStudy (anonymous):

help: how can you solve y ' cos x = sin(2x)

jimthompson5910 (jim_thompson5910):

isolate y' and use the idea that sin(2x) = 2*sin(x)*cos(x) and simplify then integrate both sides with respect to x to get y

OpenStudy (mathmale):

Is that the derivative of y? y' ? sin 2x If so, would it be true that y' = ---------- ?? cos x

OpenStudy (mathmale):

Jim and I are suggesting pretty much the same thing.

OpenStudy (mathmale):

Bet you're working on simplifying that equation yourself, but if not:\[\frac{ dy }{ dx }=\frac{ 2\sin x \cos x }{ \cos x }=?\]

OpenStudy (mathmale):

@SalvadorV? Still interested in solving this problem? Lots of other interesting problems posted on OpenStudy, no?

OpenStudy (mathmale):

Bedtime for me. Hope to see you on OpenStudy again. Buenas Noches.

OpenStudy (anonymous):

sorry, I forgot to check back in this section. Thank you very much, I then solve it as \[\frac{ dy }{ dx }=2\sin x=2\cos x\]

OpenStudy (mathmale):

Very sorry for the delays, Salvador. I've lost track of where your Laplace transform post is; can you direct me back to it by sharing the URL?

OpenStudy (anonymous):

sure, and thank you very much http://openstudy.com/study#/updates/5359aa39e4b0fb65c3cad2ec